A buffer contains 0.19 mol of propionic acid (C2H5COOH) and 0.17 mol of sodium propionate (C2H5COONa) in 1.20 L.What is the pH of the buffer after the addition of 0.02 mol of HI?
Answer #1

pH of acidic buffer = pka + log(C2H5COONa/C2H5COOH)

after addition of acid HI

pH of acidic buffer = pka + loglog(C2H5COONa- HI/C2H5COOH+ HI)

pka of CH3CH2COOH = 4.87

Number of moles of HI added is 0.02

After addition of acid HI, the pH of acidic buffer = pka + loglog(C2H5COONa- HI/C2H5COOH+ HI)

                                                                            = 4.87+ log((0.17-0.02)/(0.19-0.02)) = 4.81

After addition of acid HI, the pH of acidic buffer is 4.81